### v3

# A Context-Contrastive Inference Approach To Partial Diacritization

Abstract
Diacritization plays a pivotal role in improving readability and disambiguating the meaning of Arabic texts. Efforts have so far focused on marking every eligible character (Full Diacritization). Comparatively overlooked, Partial Diacritzation (PD) is the selection of a subset of characters to be annotated to aid comprehension where needed. Research has indicated that excessive diacritic marks can hinder skilled readers—reducing reading speed and accuracy. We conduct a behavioral experiment and show that partially marked text is often easier to read than fully marked text, and sometimes easier than plain text. In this light, we introduce Context-Contrastive Partial Diacritization ()—a novel approach to PD which integrates seamlessly with existing Arabic diacritization systems. processes each word twice, once with context and once without, and diacritizes only the characters with disparities between the two inferences. Further, we introduce novel indicators for measuring partial diacritization quality to help establish this as a machine learning task. Lastly, we introduce TD2, a Transformer-variant of an established model which offers a markedly different performance profile on our proposed indicators compared to all other known systems.111Demo: https://huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization
A Context-Contrastive Inference
Approach To Partial Diacritization
Muhammad ElNokrashy Microsoft Egypt muelnokr@microsoft.com Badr AlKhamissi EPFL Switzerland badr.alkhamissi@epfl.ch
1 Introduction
The Arabic language is central to the linguistic landscape of over 422 million speakers. It plays a pivotal role in the religious life of over a billion Muslims \@BBOPcite\@BAP\@BBNMijlad and El Younoussi (2022)\@BBCP. As in other impure abjad writing systems, the Arabic script omits from writing some phonological features, like short vowels and consonant lengthening. This can affect reading efficiency and comprehension. Readers use context from neighbouring words, the domain topic, and experience with the language structure to guess the correct pronunciation and disambiguate the meaning of the text.
The Arabic NLP community has noticeably focused on the task of Full Diacritization (FD)—the modeling of diacritic marks on every eligible character in a text (for example: \@BBOPcitep\@BAP\@BBNDarwish et al., 2017; Mubarak et al., 2019; AlKhamissi et al., 2020\@BBCP). This is especially useful in domains where ambiguities are not allowed, or where deducing the correct forms might pose challenges for non-experts. Such domains may include religious texts, literary works like poetry, or educational material.
There are benefits for human readers, like facilitating learning. However, prior research suggests that extensive diacritization can inadvertently impede skilled reading by increasing the required processing time \@BBOPcite\@BAP\@BBNTaha (2016); Ibrahim (2013); Abu-Leil et al. (2014); Midhwah and Alhawary (2020); Roman and Pavard (1987); Hermena et al. (2015)\@BBCP. Nevertheless, diacritics are important morphological markers, even when excessive, and may benefit automated systems in language modeling, machine translation (MT), part-of-speech tagging, morphological analysis, acoustic modeling for speech recognition, and text-to-speech synthesis. For an example in MT, see \@BBOPcitet\@BAP\@BBNFadel et al. (2019); Habash et al. (2016); Alqahtani et al. (2016); Diab et al. (2007)\@BBCP.
System | Sentence | |
Full Diacritization | ||
[…]
–MV, hard– | Partial Diacritization | |
[…]
1.1 Contributions
Context-Contrastive Partial Diacritization
We propose a novel method for Partial Diacritization (PD) which seamlessly utilizes existing Arabic FD systems. We exploit a statistical property of Arabic words wherein readers can guess the correct reading of most unmarked words with minimal context. To select the letters to mark, sees each word twice: (1) within its sentence context, and (2) as an isolated input with no context. The two predictions are combined to retain only those diacritics which present comparatively new information that may aid reading comprehension (Section 4).
Human Evaluation
We conduct a behavioral experiment to compare ease of reading for text with all, some, or no diacritics. Diacritics are selected via a neural model mask to simulate native partial marking. The results support prior work on the effect of different degrees of diacritcs on reading efficiency and comphrension, and is the motivation for this work (see Section 8).
Performance Indicators
We introduce a set of automatic indicators to gauge the performance and usefulness of our method on several public models, in light of the scarcity of supervised labeled partially diacritized datasets (see Section 5).
Transformer D2
And last, we present the TD2 model, a Transformer variant of D2 \@BBOPcite\@BAP\@BBNAlKhamissi et al. (2020)\@BBCP, which shows markedly improved Partial DER performance at compared to , even while marking a larger percentage of text at compared to for D2 (see Section 6.1).
Glyph | Name | Type | BW | IPA |
[…]
2 Motivation
Previous research underscores the substantial influence of diacritization on the reading process of Arabic text. Skilled readers exhibit a tendency to read highly diacritized text at a slower pace compared to undiacritized. The effect on reading speed and accuracy has been substantiated across numerous studies involving diverse age groups and linguistic backgrounds \@BBOPcite\@BAP\@BBNTaha (2016); Ibrahim (2013); Abu-Leil et al. (2014); Midhwah and Alhawary (2020); Roman and Pavard (1987); Hermena et al. (2015); Hallberg (2022)\@BBCP. In addition, studies monitoring eye movements during reading link the increased reading times for diacritized text to an increase in fixation frequency and duration—a key indicator of the word identification process \@BBOPcite\@BAP\@BBNRoman and Pavard (1987); Hermena et al. (2015)\@BBCP. Partial Diacritization may thus lead to a better reading experience for a wide range of readers, including in cases with dyslexia or visual impairments.
3 Arabic Preliminaries
Diacritics in Modern Arabic Orthography
The inventory of diacritics used in modern typeset Arabic comprises at least four functional groups. First are ḥarakāt and sukūn, which indicate vowel phonemes or their absence. Then tanwīn and shaddah indicate case-inflection morphemes and consonant lengthening. There are other marks, e.g. for recitation features in the Qur’an. Only the marks in Table 2, part of modern Arabic orthography, are discussed here. Ḥarakāt are vowel diacritics that provide vowel information. Sukūn indicates the absence of a vowel, which indicates consonant clusters or diphtongs. Tanwīn diacritics indicate the phonemic pair of a vowel and the consonant /n/ at the end of words, serving case inflection purposes. These diacritics play essential roles in Arabic orthography and pronunciation \@BBOPcite\@BAP\@BBNHallberg (2022)\@BBCP.
Modes of Diacritization.
Since diacritics are optional and their usage can vary widely—practical patterns have emerged. \@BBOPcitet\@BAP\@BBNHallberg (2022)\@BBCP identifies seven modes of diacritization. These modes can be ordered based on the quantity of diacritics used, ranging from no diacritization to complete diacritization. Deeper levels, like complete diacritization, are less frequently used than shallower levels (no diacritization, and so on). Examples in Figure 2.
4 Methods
4.1 Full Diacritization Models
We utilize deep neural sequence models for character diacritization. We test the pretrained models D2, Shakkelha and Shakkala \@BBOPcite\@BAP\@BBNAlKhamissi et al. (2020); Fadel et al. (2019); Barqawi (2017)\@BBCP. Further, we introduce the TD2 model: a Transformer variant of the LSTM-based D2. A description of TD2 and the D2 architecture can be found at Section 6.1.
4.2 What Context Do Models See?
Context extraction.
Let contextualize from . Then the word and letter context extractors are functions which use the surrounding words and letters around positions and .
Let the sub-sequence contain all the words in a segment of length s.t. as used in training D2/TD2.
[…]
Thus refers to a segment containing the word , while refers to the intra-word context for letter in word .
Contextual Prediction
sees the full segment.
Single Word Prediction
sees only the current word222Sentences are split on spaces. Words are not tokenized. at position . Note the instead of .
Thus the two modes are defined as follows: is single-word application, while uses sentence context. Both use the same model .
4.3 Context-Contrastive Partial Diacritization
The function takes each word in isolation. If its prediction is correct, we assume that this reading is common and easy to guess by a human reader. By contrast, utilizes word context within a sentence. If a word has multiple possible readings which the sentence disambiguates, we expect with sentence context to out-perform .
Following from these premises, we propose the following method to combine both sources of information to diacritize text partially, with justification.
CCPD: Algorithm
Using and predictions for a word and letter , the system assigns or omits a diacritic according to :
In hard or disagreement mode, the letters where inferences with and without context agree are left bare (regardless of the correctness of the prediction; unknown during inference). Otherwise, the contextual prediction is returned.
In soft or confidence mode, only the letters where the logit exceeds the logit by a margin receive a diacritical mark.
5 Performance Indicators
To measure performance on the partial diacritization task, we propose four indicators as approximate gauges for model alignment with human expectation. These indicators aim to address the challenge of the limited availability of high-quality partially diacritized test datasets, providing valuable insights into the model’s capabilities.
Indicators may use the marked () and unmarked () subsets of the corpus ().
selected | (8) | ||||
unmarked |
5.1 SR: Selection Rate
SR is the proportion of characters assigned a diacritic by Function (LABEL:eq:pdd-assign). Literature shows partial diacritization hovers around – coverage in some professionally published books \@BBOPcitep\@BAP\@BBN(Hallberg, 2022)\@BBCP, while other research suggests that deliberate partial diacritization by native speakers results in a rate around – \@BBOPcitep\@BAP\@BBN(Esmail et al., 2022)\@BBCP.
5.2 Scoped Diacritic Error Rates DER(f, D)
[…]
5.2.1 P-DER: Partial Diacritic Error Rate
[…]
5.2.2 B-DER: Basic Diacritic Error Rate
By the intuition in Section 4.3 regarding the non-contextual predictions of , we calculate the DER on the whole corpus and use B-DER as a proxy for human error on plain unmarked text.
5.3 RE-DER: Reader DER
We combine the Partial DER and Basic DER into one general measure of the total error in a partially diacritized text. The Basic DER indicator is used as a proxy which may be taken as an upper bound for mistakes made by an experienced reader. Formally, this indicator incorporates the error of the non-contextual model for the unmarked text, and of the contextual-model for the marked text (because the system explicitly annotates it). Thus RE-DER (Reader DER) is:
5.4 SU: Signal Utilization
How well does utilize its marked subset to disambiguate the text? We gauge how informative the added diacritics are by measuring the change in DER within the marked subset of text. SUspans where positive values denote improvement over no annotation. SU is defined as:
E.g. an SU=50% shows that half of the produced annotations clarify ambiguities (lower RE-DER).
System | SR | SU | P-DER | RE-DER | B-DER | DER | WER | |
Reference Range | ||||||||
\@BBOPcitep\@BAP\@BBNBarqawi, 2017\@BBCP | 8.6 | 70.9 | 17.9 | †3.6 | 9.7 | 3.58 | 11.19 | |
\@BBOPcitep\@BAP\@BBNFadel et al., 2019\@BBCP (big)* | 8.9 | 86.5 | 7.8 | †1.6 | 9.3 | 1.60 | 5.08 | |
\@BBOPcitep\@BAP\@BBNAlKhamissi et al., 2020\@BBCP - D2 | 6.5 | 81.5 | 11.2 | †1.8 | 7.1 | 1.85 | 5.53 | |
—Ours— | D2 (SP, hard) | 6.8 | 75.0 | 15.0 | †2.0 | 7.1 | 2.00 | 6.42 |
D2 (SP, soft ) | 5.9 | 47.5 | 3.1 | 4.3 | ||||
D2 (SP, soft ) | 11.5 | 37.4 | 4.7 | 2.8 | ||||
D2 (SP, soft ) | 15.2 | 31.6 | 5.1 | 2.3 | ||||
D2 (SP, soft ) | 21.8 | 23.4 | 4.9 | 2.0 | ||||
TD2 (MV, hard) | 24.6 | 91.5 | 5.5 | †2.4 | 25.0 | 2.44 | 7.68 |
6 Experimental Modeling
6.1 TD2: Transformer D2
The TD2 model is a Transformer adaptation of D2 \@BBOPcitep\@BAP\@BBN(AlKhamissi et al., 2020)\@BBCP. It comprises two encoders of 2 layers each, for tokens and characters. Both models use the same architecture. The feature and intermediate widths are (768, 2304). The token model is initialized from the pretrained weights of layers 1-2 of CAMeL-Lab/bert-base-arabic-camelbert-mix-ner \@BBOPcitep\@BAP\@BBN(Inoue et al., 2021)\@BBCP as provided by HuggingFace333HuggingFace: \@BBOPcitet\@BAP\@BBNWolf et al. (2020)\@BBCP. The character model is initialized from layers 3-4 of the same pretrained weights.
6.2 SP: Single-Pass Inference
For recurrent (non-Transformer) models like D2, we also test inference without segmentation and majority voting. Each sentence is passed in full the model, once. This can save inference time (by avoiding overlapping windows), but sacrifices some reliability (as the model had been trained on smaller window sizes).
6.3 Other Models
7 Results
7.1 Full Diacritization Performance
Following prior work, we report the Diacritic Error Rate (DER) and Word Error Rate (WER) including case-endings which are located at the word’s end, usually determined by the word’s syntactic role. Predicting these diacritics is more challenging compared to core-word diacritics, which specify lexical properties and lie elsewhere within the word.
7.2 Partial Diacritization Performance
Intuition
Overall it is desirable to observe: (1) a natural SR value close to native human annotation rates; (2) a low B-DER close to the SR, signifying a capable model which allows a clean partial annotation of only the hardest letters, (3) a low P-DER to signify high accuracy in the committed diacritics, and (4) a low RE-DER which gauges the overall expected reading experience given the partial annotation and the expected guessing error. Overall, this corresponds to a balance between a high SU, a natural SR, and a low P-DER, roughly ordered.
Model results
Notice that the recurrent D2 is conservative in its selection with SR at 6.46% versus TD2’s 24.61%.
TD2 under-performs in non-contextual mode with B-DER at 25% compared to 7.1% by D2. Counter-intuitively, this leads it to a higher SU at 91.5% as its contextual mode corrects 22.5 of every 25 errors its non-contextual mode commits.
Both systems agree that much of the text (at least 75%) is reasonably guessable by the reader without annotation, by taking B-DER as an upper-bound on human guess error.
We may claim that between 75% and 93% of characters require no diacritization to disambiguate (the extremes of the two systems).
Both D2 and TD2 correct most of their B-DER error via , implying that most letters which had errors in non-contextual mode were accurately selected by and correctly predicted by , leading to an RE-DER < B-DER.
Soft marking
SR can be tweaked in soft selection mode. Let’s analyze the model with SR=15.2% (Table 3) which provides a low P-DER at 5.1% (few diacritics added are erroneous) and a low RE-DER at 2.3% (around 4.8 out of 7.1% expected errors are corrected via the added diacritics). Notice the low SU (utilization) however, as only 31.6% of the annotated letters contribute to the drop in error.
Assuming that the selected letters are the hardest to guess for a reader, this suggests that D2-SP models can disambiguate the text by 39.4% to 71.8% (relative improvement in RE-DER/B-DER) by marking only 5.9% to 21.8% of it.
While RE-DER improves, P-DER and SU worsen, suggesting that increased annotation may give only marginal improvement in overall readability if the final model output is not perfectly clean.
8 Behavioral Experiment on Partial Diacritization
We start with the hypothesis that reading partially diacritized text may be easier than reading fully diacritized text, while providing some benefit in disambiguation or ease-of-reading.
To test this hypothesis, we conduct a behavioral experiment using machine-predicted partial diacritization masks. The diacritical marks themselves are the true labels. This is to ensure that the presented text is accurate, even if the model output is sub-optimal. The question is then: Does using and select useful letters to be diacritized? See the results in Figure 4.
8.1 Behavioral Experiment Setup
Demographics
We utilize the PsyToolkit online platform444PsyToolkit: \@BBOPcitet\@BAP\@BBNStoet (2010, 2017)\@BBCP to conduct a behavioral experiment aimed at assessing the impact of different levels of diacritization (modes) on reading speed and accuracy. Data was gathered from a group of 15 participants, covering various age groups and native speakers of different Arabic dialects. The majority of participants (10) fell within the 25–40 age group; 2 were below 25, and 3 were above 40. All participants spoke at least 2 languages, while 6 spoke 3 or more. Eight participants spoke either Gulf (4) or Maghrebi (4) Arabic natively, and 7 spoke Egyptian (2), Levantine (2), or Sudanese (2). One did not report a native dialect. All participants reported being native speakers of Arabic.
Dataset
To measure the impact of diacritics, we select 30 sentences from various domains in the WikiNews testset \@BBOPcitep\@BAP\@BBN(Darwish et al., 2017)\@BBCP. Each sentence is presented in three variants with Zero/Partial/Full diacritics. The Partial variant is produced using our algorithm and the D2-SP model in selection mode to mask out easy or guessable ground truth diacritics.
Data Splits
The data is split into 3 buckets of 10 sentences each. We create 3 test sets such that the buckets are rotated and each appears exactly once in any mode (Zero/Partial/Full). For example, the first bucket would appear in the test sets A, B, C in its Full, Zero, and Partial variants. This is done to ensure that any participant sees a sentence exactly once, to avoid biasing their rating by repeated exposure. This ensures also that each sentence is seen equal times in each mode.
Timing and Scoring
Participants are shown each sentence for a few seconds proportional to the word count of the sentence: . Then the participants are prompted to rate their comprehension by a score from 1 to 5, where 1 indicates difficulty and 5 indicates ease in understanding the sentence. The design intention is to measure the reading experience of a native speaker when scanning a non-technical text in Modern Standard Arabic.
8.2 Findings
Figure 4 illustrates the per-mode average self-reported reading comprehension scores, normalized by the score of the no diacritization (Zero) mode. The results indicate that Full diacritization hinders reading accuracy when participants are provided with a limited amount of time to read the text, aligning with findings discussed in Section 2. In contrast, the Partial diacritization mode sometimes enhances reading comprehension performance compared to the Zero mode.
The data from this behavioral experiment will be available on the paper’s GituHub repository555GitHub: munael/arabic-partial-diacritization.
9 Related Work
The literature has dealth with both types of diacritization: Full (FD) and Partial (PD). The majority has focused on FD. In PD, a distinct approach that seeks to augment reading comprehension by incorporating only the minimal requisite diacritics, we note a few works focus \@BBOPcitep\@BAP\@BBN(Almanea, 2021; Mijlad and El Younoussi, 2022)\@BBCP. Other works take a morphological analysis approach \@BBOPcitep\@BAP\@BBN(Obeid et al., 2022, 2020; Alqahtani et al., 2016; Shahrour et al., 2015; Habash and Rambow, 2007)\@BBCP.
9.1 Full Diacritization
Non-Neural Methods
This class of methods combines linguistic rules with non-neural modelling techniques, such as Hidden Markov Models (HMMs) or Support Vector Machines (SVMs), which were widely employed over a decade ago. For example, \@BBOPcitet\@BAP\@BBNElshafei et al. (2006)\@BBCP utilize an HMM to predict diacritization based on bigram and trigram distributions. \@BBOPcitet\@BAP\@BBNBebah et al. (2014)\@BBCP extract morphological information and utilize HMM modeling for vowelization, considering word frequency distribution. \@BBOPcite\@BAP\@BBNShaalan et al. (2009)\@BBCP combine lexicon retrieval, bigram modeling, and SVM for POS tagging, addressing inflectional characteristics in Arabic text. \@BBOPcitet\@BAP\@BBNDarwish et al. (2017)\@BBCP operate diacritization in two phases, inferring internal vowels using bigrams and handling case endings via an SVM ranking model and heuristics. \@BBOPcitet\@BAP\@BBNSaid et al. (2013)\@BBCP follow a sequence-based approach, involving autocorrection, tokenization, morphological feature extraction, HMM-based POS tagging, and statistical modeling for handling OOV terms. \@BBOPcitet\@BAP\@BBNZitouni and Sarikaya (2009)\@BBCP approaches the problem using maximum entropy models.
Neural Methods
Recent literature has focused more on neural-based systems. \@BBOPcitet\@BAP\@BBNBelinkov and Glass (2015)\@BBCP were the first to show that recurrent neural models are suitable candidates to learn the task entirely from data without resorting to manually engineered features such as morphological analyzers and POS taggers. \@BBOPcitet\@BAP\@BBNAlKhamissi et al. (2020)\@BBCP used a hierarchical, BiLSTM-based model that operates on words and characters separately with a cross-level attention connecting the two—enabling SOTA task performance and faster training and inference compared to traditional models. Many prior works have used recurrent-based models for the FD task \@BBOPcite\@BAP\@BBNAl-Thubaity et al. (2020); Darwish et al. (2020); Gheith Abandah (2020); Fadel et al. (2019); Boudchiche et al. (2017); Moumen et al. (2018)\@BBCP, others used the Transformer architecture \@BBOPcite\@BAP\@BBNMubarak et al. (2019)\@BBCP, while others used ConvNet architectures \@BBOPcite\@BAP\@BBNAlqahtani et al. (2019b)\@BBCP.
Hybrid Methods
9.2 Partial Diacritization
Similar to this work, previous research has explored the selective diacritization of Arabic text. \@BBOPcitet\@BAP\@BBNAlnefaie and Azmi (2017)\@BBCP harnessed the output of the MADAMIRA morphological analyzer \@BBOPcite\@BAP\@BBNPasha et al. (2014)\@BBCP and leveraged WordNet to generate word candidates for diacritics. This work focused on resolving word ambiguity through statistical and contextual similarity approaches to enhance diacritization effectiveness. \@BBOPcitet\@BAP\@BBNAlqahtani et al. (2019a)\@BBCP focused on selective homograph disambiguation, proposing methods to automatically identify and mark a subset of words for diacritic restoration. Evaluation of various strategies for ambiguous word selection revealed promising results in downstream applications, such as neural machine translation, part-of-speech tagging, and semantic textual similarity, demonstrating that partial diacritization effectively strikes a balance between homograph disambiguation and mitigating sparsity effects. \@BBOPcitet\@BAP\@BBNEsmail et al. (2022)\@BBCP employ two neural networks to predict partial diacritics—one considering the entire sentence and the other considering the text read so far. Partial diacritization decisions are made based on disagreements between the two networks, favoring the prediction conditioned on the whole sentence.
10 Discussion
In this work, we have presented a novel approach for partial diacritization using context-contrastive prediction. The task is motivated by a large body of literature on the impact of diacritization coverage on the reading process of Arabic text, and the scarcity of research into systems to automate it.
Context-Contrastive Prediction
By considering contextual information alongside non-contextual predictions, our approach builds on the role of diacritics in text, particularly in the context of the Arabic script as used by humans where accuracy is not the only consideration. The result is a simple, efficient, and configurable method which can be integrated with any existing diacritization system. It opens doors to enhanced diacritization accuracy and selection. The indicators we introduce provide valuable signals for evaluating such advancements.
The Effect of Diacritics on Reading
A significant motivation for our work rests upon prior research that into the substantial influence of diacritization on the reading process of Arabic text. Multiple studies have consistently shown that skilled readers tend to read extensively diacritized text at a slower pace than undiacritized text. This phenomenon, supported by research involving diverse age groups and linguistic backgrounds, has significant implications for reading accuracy and comprehension \@BBOPcitep\@BAP\@BBN(Taha, 2016; Ibrahim, 2013; Abu-Leil et al., 2014; Midhwah and Alhawary, 2020; Roman and Pavard, 1987; Hermena et al., 2015; Hallberg, 2022)\@BBCP. Other studies also suggest that diacritization not only affects reading speed but also directly impacts reading accuracy \@BBOPcitep\@BAP\@BBN(Roman and Pavard, 1987; Hermena et al., 2015)\@BBCP. In this work, we complement this large body of research by conducting a behavioral experiment that utilized our approach to partially diacritize news headlines from the WikiNews test set. Our results show that partially diacritized text is easier to read than fully diacritized text, and in some cases can lead to better understanding than no diacritization at all.
Implications
An exciting potential for this work is optimizing the reading experience for different groups of readers with different needed levels of diacritization, including people with dyslexia and visual impairments. Since diacritization plays an important role in disambiguating the meaning of the text, it is crucial to intelligently select the ones that aid readability and comprehension of Arabic text without excessive marking; making it accessible and accommodating to a broader audience. This aligns with the broader goals of inclusivity and accessibility in NLP applications.
Future Directions
Further research is needed to gather high-quality benchmark data for partial diacritization, to enable more traditional and direct performance metrics alongside the indicators we propose. One promising direction is evaluating the method on other languages which utilize diacritical marks in their orthographies.
11 Conclusion
In conclusion, our approach for partial diacritization using context-contrastive prediction contributes to the field of diacritization in an area with far-reaching implications for enhancing the reading experience of Arabic text. Grounded in previous research on the influence of diacritization on reading and further supported by behavioral experiments conducted in this work—our approach paves the way for advancements that benefit readers of diverse backgrounds and abilities. Since our approach integrates seamlessly with existing Arabic diacritization systems, it can be used by models trained on domain-specific text as well. We also propose a battery of performance indicators to gauge the competence of partial diacritization systems using fully-diacritized test data to mitigate the lack of publicly available benchmarks. Finally, we introduce TD2—a Transformer adaptation of the D2 model which offers a different performance profile as shown by our proposed indicators. As Arabic NLP continues to evolve, our approach serves as a promising direction for enhancing diacritization and its impact on text accessibility and comprehension.
Limitations
The method presented in this work is tested on a single human language. While we believe it should be able to generalize to other languages that need similar diacritics restoration, its utility needs further observation. In particular, it may be the case that partial diacritization patterns in languages besides Arabic are different in such a way that the premise of this work no longer holds (wherein ease of guessing is regarded as a major factor in omitting a diacritical mark).
In addition, this method is deliberately kept simple, which builds on an over-simplified view of the human task. Readers guess diacritics when reading, but they may do so while incorporating context via some simple or fast mechanisms. Whether that is at a deep level equivalent to guessing the reading of a single word in isolation remains to be seen. Our method uses non-contextual application, , as a proxy for this process, which is likely relatively close in performance to the same mechanism in a human. Nevertheless, we make no claim that this is indeed the natural mechanism, nor that the proxy exhibits identical performance distribution to the natural mechanism.
Appendix
Appendix A Experimenting with [Transformer] D2
A.1 Datasets
In this work, we train on data from the Tashkeela \@BBOPcitep\@BAP\@BBN(Fadel et al., 2019)\@BBCP and Ashaar \@BBOPcitep\@BAP\@BBN(Alyafeai et al., 2023)\@BBCP corpora. We test on the Tashkeela testset. The Tashkeela corpus has been collected mostly from Islamic classical books \@BBOPcite\@BAP\@BBNZerrouki and Balla (2017)\@BBCP and contains mostly classical Arabic sentences. Ashaar is a corpus of Arabic poetry verses covering poems from different eras. Table 4 shows the number of tokens for the Tashkeela and Ashaar666The reported numbers reflect dataset cleaning to keep only Arabic letters and diacritics. datasets.
Train | Dev | Test | |
Tashkeela | |||
Ashaar |
A.2 Majority Voting using a Sliding Window
Following prior work that utilizes an overlapping context window approach with a voting mechanism to enhance diacritic prediction for individual characters \@BBOPcitep\@BAP\@BBN(Mubarak et al., 2019)\@BBCP, we segment each input sentence into multiple overlapping windows. We present each segment to the model separately. This approach has proven effective, as localized context often contains enough information for accurate predictions. We use similar segmentation parameters to \@BBOPcitet\@BAP\@BBNAlKhamissi et al. (2020)\@BBCP
Training
For the training and validation sets we use a window of words and a stride of .
Inference
The same character may appear in different contexts, and therefore potentially result in different diacritized forms. We implement a popularity voting mechanism to narrow down the prediction. When a tie arises, we randomly select one of the top options. Testing is done with a sliding window of words at a stride of .
A.3 Datapath of D2, TD2
The token model’s output is averaged per word to result in exactly one feature vector (instead of one for each sub-word). This aggregated vector is concatenated with the character embedding of each character in the word and down-projected to the model feature size, to get character input . The character encoder transforms the inputs into character feature vectors which are passed to the final classifier.
Training
The full model is tuned using the AdamW optimizer \@BBOPcitep\@BAP\@BBN(Loshchilov and Hutter, 2019)\@BBCP with dropout and LR, and follows a linear schedule of warmup steps and total training steps. The best checkpoint at step intervals is picked by the combined Tashkeela and Ashaar dev sets.
Appendix B Demo and Examples
We developed an online Demo on Huggingface777huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization which allows choice between Full Diacritization, and Partial Diacritization with Hard and Soft modes. See Figures 5(a) and 5(b). The demo supports the D2/TD2 models (trained on Tashkeela and Tashkeela/Ashaar). Since they focus mostly on classical Arabic text, users are advised not to anticipate optimal performance when applying this model to Modern Standard Arabic (MSA) or informal Arabic dialects.
B.1 More Partial Diacritization Examples
The following examples are lines taken from a poem by Rami Mohamed (2016) called “We Will Stay Here”. We applied the models as indicated in the System column. The input text included no diacritics at all. See Table 5.
# | System | Text/Output | Error Count |
[…]
Translation | We will Stay Here .. That the Pain may one day Cease | ||
[…]
Translation | Let Us all Rise .. With Healing and Writing | ||
[…]
Translation | My Joys and Screams .. Through Deafness Nearly Heard | ||
[…]